What is the critical speed(RPM) of the ball mill, if the diameter of the ball mill=500 mm and diameter of the ball 80 mm


A) 65


B) 75


C) 45


D) 55


E) 35


1
Expert's answer
2021-08-18T09:36:01-0400

"N_c= \frac{1}{2 \pi}[\frac{g}{R-r}]^{0.5}\\
N_c= \frac{1}{2 \pi}[\frac{9.8*10^3}{500-80}]^{0.5}\\
=\frac{1}{2\pi }\left(\frac{9.8\cdot \:10^3}{500-80}\right)^{\frac{1}{2}}\\
=\frac{5\cdot \:19.6^{\frac{1}{2}}}{21^{\frac{1}{2}}}\\
=\frac{1}{2\pi }\cdot \frac{5\cdot \:19.6^{\frac{1}{2}}}{21^{\frac{1}{2}}}\\
=\frac{22.13594\dots }{28.79317\dots }\\
=0.76879*100\\
=76.9"